Please Help,
What is the product of 2Radical 7 + 3 radical 5 and is conjugate?
also
Solve the following equations:
a) 4 + radical x+1 = 5
b) 2/7 = 5/3x-1
c) c-7/c+2 = 1/4
Thank you for your help...

Answer by stanbon(75887)

What is the product of 2Radical 7 + 3 radical 5 and is conjugate?
(2sqrt7 + 3sqrt5)(2sqrt7 - 3sqrt5)
= 4*7 -9*5 = 28-45 = -17
-----------------------------
also
Solve the following equations:
a) 4 + radical x+1 = 5
sqrt(x+1) = 1
Square both sides to get:
x+1=1
x=0
---------------
b) 2/7 = 5/3x-1
Cross-multiply to get:
2(3x-1) = 35
6x-2 = 35
6x = 37
x = 37/6
----------------
c) c-7/c+2 = 1/4
Cross-multiply to get:
4(c-7) = c+2
4c-27 = c+2
3c = 29
c = 29/3
==============
Cheers,
Stan H.